OpenStudy (anonymous):

Solve 3^x – 4 = 7^x + 9

OpenStudy (anonymous):

x ≈ –25.86 x ≈ –2.59 x ≈ –0.39 x ≈ –0.04

OpenStudy (phi):

if this is \[ 3^{x-4}= 7^{x+9} \] can you take the log of both sides ?

OpenStudy (anonymous):

How would i go about doing that?

OpenStudy (phi):

you write log next to each term

OpenStudy (anonymous):

Like log3^x-4 = log7^x+9?

OpenStudy (phi):

yes, but put in parens log (3^(x-4) ) = log(7^(x+9) ) which looks nicer if you use the equation editor (see button on the bottom left of where you type in) \[ \log\left(3^{x-4}\right) = \log\left(7^{x+9}\right) \]

OpenStudy (phi):

now use this property \[ \log(a^b) = b\log(a) \] to rewrite both sides.

OpenStudy (anonymous):

What is a & b?

OpenStudy (phi):

you do pattern matching. match the rule to your expression.

OpenStudy (anonymous):

log(3^x4) = ()7log(x+9)?

OpenStudy (phi):

Here is an example of how to interpret the rule \[ \log(a^b) = b\log(a) \] if you have the problem \[ \log(y^3) \] you match the y to the a (both are the "base") and match 3 to the b (both are "exponents") once you figure out that the 3 matches the b, you use the 2nd part \[ b\log(a) \] to re-write it as \[ 3 \log(y) \]

OpenStudy (phi):

in your case, what matches to the a ? what matches the b? (what is the exponent)

OpenStudy (phi):

use the rule on the left side of \[ \log\left(3^{x-4}\right) = \log\left(7^{x+9}\right) \] can you match \[ \log\left(3^{x-4}\right) \text{ to } \log(a^b) \]

OpenStudy (phi):

what matches to the a ? what matches the b? (what is the exponent)